OpenStudy (anonymous):

Mixture and Uniform Motion Problem (Help Please) A clothing manufacturer purchased 65 yd of cotton and 90 yd of wool for a total cost of $2,000. Another purchase, at the same prices, included 90 yd of cotton and 25 yd of wool for a total cost of $1,275. Find the cost per yard of the cotton and of the wool.

OpenStudy (caozeyuan):

do you have your equations ready?

OpenStudy (anonymous):

No I don't understand how to do this problem.

OpenStudy (caozeyuan):

say one yard of cotton is c dollar, one yard of wool is w dollar, what are the components of the 2000 dollar? answer is a combination of c and w

OpenStudy (anonymous):

cotton(c dollar)+ wool(w dollar)=2000

OpenStudy (caozeyuan):

one strike! did you forgot something?

OpenStudy (anonymous):

cotton(c dollar)+ wool(w dollar)=1275?

OpenStudy (caozeyuan):

two strikes

OpenStudy (caozeyuan):

you forgot the price

OpenStudy (caozeyuan):

nope, the area

OpenStudy (anonymous):

I'm so confused

OpenStudy (caozeyuan):

so 65c+90w=2000, whats the other one

OpenStudy (anonymous):

90c+25w=1275

OpenStudy (caozeyuan):

great, now we can say 324c+90w=4590, understood

OpenStudy (anonymous):

yes

OpenStudy (caozeyuan):

so 259c=2590,c=10

OpenStudy (anonymous):

Did u divide?

OpenStudy (anonymous):

Where did u get 259?

OpenStudy (anonymous):

@caozeyuan

OpenStudy (caozeyuan):

4590-2000=2590

OpenStudy (anonymous):

Ok where do I go from here?

OpenStudy (caozeyuan):

throw c back in to get w out

OpenStudy (anonymous):

so plug it in

OpenStudy (anonymous):

Wool=15 @caozeyuan

OpenStudy (caozeyuan):

great you did it( insert applause here)

OpenStudy (anonymous):

Thanks. So the cost per wool is $15?

OpenStudy (anonymous):

@caozeyuan

OpenStudy (anonymous):

And $10 is the cost per yard for cotton?

OpenStudy (caozeyuan):

yes

OpenStudy (anonymous):

that's it?